OpenStudy (anonymous):

Find the point on the line 6x+y=9 that is closest to the point (−2,8)

OpenStudy (anonymous):

Let P(x,y) is the point on 6x + y = 9 such as ( x, -6x +9 ) Find the distance to ( -2, 8) Take derivative!

OpenStudy (anonymous):

How do i take the derivative of two unknowns?

OpenStudy (anonymous):

You have only x is variable!

OpenStudy (anonymous):

so 6?

OpenStudy (anonymous):

How do you get 6?

OpenStudy (anonymous):

y = 9-6x y = 2(3-2x) ?

OpenStudy (anonymous):

It's okay. I got the answer, so i'm pretty sure it's fine :)